angelique27:

A vending machine sells chips at $0.55 and candy at $0.75 . Last month, the vending machine yielded $189.00 with the sale of 280 items

darkknight:

f for chips c for candy .55f + .75c = 189 f + c = 280

darkknight:

do u understand how i got those 2 eqns?

angelique27:

no

angelique27:

whats the answer

darkknight:

the cost function is one eqn, the amount of items total is another

darkknight:

i am trying to help you get the answer

angelique27:

ok thanks

angelique27:

so what do i do to find the answer

darkknight:

i gave u 2 eqns, first do u get how i got them?

angelique27:

yes

darkknight:

nice now .55f + .75c = 189 f + c = 280 c = 280-f now we can use substitution

darkknight:

.55f + .75(280-f) = 189 i plugged in the value for c so that our eqn is in terms of f

darkknight:

now i would like you to solve for f

angelique27:

you multiply 75 with 280 right?

darkknight:

.75

darkknight:

you have to distribute within the parenthesis

angelique27:

$210.55=189 right?

darkknight:

?

darkknight:

check your math, you should be solving for f

angelique27:

i did

angelique27:

nvm this is too hard

darkknight:

no it isn't u got this .55f + .75(280-f) = 189 you have to solve for f, just distribute and combine like terms

ZombieLover:

Close when done :)